Equity value in a levered firm.  Air Seattle has an annual EBIT of ​$1 comma 200 comma 000​, and the WACC in the unlevered firm is 17 %. The current tax rate is 35​%. Air Seattle will have the same EBIT forever. If the company sells debt for $ 2 comma 200 comma 000 with a cost of debt of 23​%, what is the value of equity in the unlevered firm and in the levered​ firm? What is the value of debt in the levered​ firm? What is the​ government's value in the unlevered firm and in the levered​ firm?

If Air​ Seattle's WACC is 17​% and it will have the same EBIT​ forever, what is the current value of Air​ Seattle?
Answer #1

Unlevered Firm:

Value of Unlevered Firm = Value of Equity

PV of Cash Flow = EBIT / WACC = $1,200,000 / 0.17 = $7,058,823.53

Value of Equity = PV of cash flow * (1 - t) = $7,058,823.53 * (1 - 0.35) = $4,588,235.29

Levered Firm:

Value of levered firm = Value of Unlevered Firm + [Debt * Tc]

= $4,588,235.29 + [$2,200,000 * 0.35] = $4,588,235.29 + $770,000 = $5,358,235.29

Value of Equity = Value of Levered Firm - Value of Debt

= $5,358,235.29 - $2,200,000 = $3,158,235.29

Government's Value:

Unlevered Firm = PV of cash flows * Tax Rate = $7,058,823.53 * 0.35 = $2,470,588.24

Levered Firm = [PV of cash flows - Amount of Debt] * Tax Rate

= [$7,058,823.53 - $2,200,000] * 0.35 = $4,858,823.53 * 0.35 = $1,700,588.24